The length of a rectangle is 1 foot less than three times the width. The area of the rectangle is 80 square feet. Find the length and width of the rectangle. Use the quadratic equation to solve.
Answer by TimothyLamb(4379)

L = 3w - 1
Lw = 80
---
Lw = 80
(3w - 1)w = 80
3ww - w - 80 = 0
3w^@ - w - 80 = 0
---
the above quadratic equation is in standard form, with a=3, b=-1, and c=-80
---
to solve the quadratic equation, by using the quadratic formula, plug this:
3 -1 -80
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two solutions), of the quadratic are:
w = 5.33333333 = 5+1/3 = 16/3
w = -5
---
the negative root doesn't make sense for width, so use the positive root:
---
answer:
w = 5.333 ft
L = 15 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php